The number of neutrons in an atom of 90Th230 is

Grade:Upto college level

2 Answers

Apoorva Arora IIT Roorkee
9 years ago
_{90}^{230}Th means mass number A= 230 and atomic number Z= 90.
We know that A=Z+N so N= 140.
Thanks and Regards
Apoorva Arora
IIT Roorkee
askIITians Faculty
Rohan kumar
2 years ago
_{90}^{230}Th means mass number A= 230 and atomic number Z= 90.
 
We know that A= Z + N =  230
                            = N = 230 - Z
                           = N = 230 - 90
                          = N =  140.
 Thanks.